# Representing Sugihara monoids

### v2

@@ Abstract
We show that all Sugihara monoids can be represented as algebras of binary relations, with the monoid operation given by relational composition. Moreover, the binary relations are weakening relations. The first step is to obtain an explicit relational representation of all finite odd Sugihara chains. Our construction mimics that of Maddux (2010), where a relational representation of the finite even Sugihara chains is given. We define the class of representable Sugihara monoids as those which can be represented as so-called direct reducts of distributive involutive FL-algebras of binary relations. We then show that the class of representable distributive involutive FL-algebras is closed under ultraproducts. This fact is used to demonstrate that the two infinite Sugihara monoids that generate the quasivariety are also representable. From this it follows that all Sugihara monoids are representable.
@@ 1 Introduction
For the odd Sugihara chain $\mathbf{S}_{2n+3}$ (where $n\in\mathbb{Z}^{+}$), we will take two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$, with a map $\alpha$ swapping the occurrence of each element from one copy to the other. This map $\alpha$ is an order isomorphism of the poset consisting of two chains. The algebra $\mathbf{S}_{2n+3}$ is then embedded in the upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$, where this time $\mathbf{X}$ consists of two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$. The key step is that $\alpha$ is used to define the order-reversing involution in a way that ensures the algebra will be an odd Sugihara monoid.
@@
The map $\alpha$ also plays a crucial role in the representation of distributive involutive FL-algebras. These algebras have two order-reversing maps $\sim$ and -- (see Sections 2.1 and 2.4). If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic. The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7). Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
@@ 2.1 Sugihara monoids and DInFL-algebras
If $n$ is odd then the monoid identity is given by $1=a_{0}$, and if $n$ is even, then the monoid identity is $1=a_{1}$. This gives us algebras $\mathbf{S}_{n}=\langle S_{n},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$. (We remark that when we talk of a Sugihara chain, we mean an algebra that is a Sugihara monoid and is linearly ordered. This is different to the terminology used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
@@ 2.3 Constructing InFL-algebras of relations
Before we give some examples, we need the following definition. It is an adaptation of [14, Definition 8].
@@
Consider the two-element Sugihara chain $\mathbf{S}_{2}$. The underlying set is given by $S_{2}=\{a_{-1},a_{1}\}$ where $a_{1}$ is the identity of the monoid operation. Consider the one-element poset with $X=\{x\}$ and order ${\leqslant}=\mathrm{id}_{X}$. Let $E=X^{2}$ and $\alpha=\mathrm{id}_{X}$. Then $\left\langle\mathsf{Up}\left(\left\langle X^{2},\leqslant\right\rangle\right),\cap,\cup,\circ,\leqslant,\sim,-\right\rangle$ is a two-element distributive InFL-algebra whose direct reduct is isomorphic to $\mathbf{S}_{2}$.
@@ 2.4 Representable DInFL-algebras
The following theorem relies on part (iii) of Theorem 2.12 and is used to representable DInFL-algebras.
@@
[3, Theorem 4.4] $\mathbb{P}\left(\mathsf{FDInFL}\right)=\mathsf{EDInFL}$
@@
In Theorem 4.11 we will show that the class of representable DInFL-algebras is closed under ultraproducts. That will be a key step in showing that all Sugihara monoids can be represented as algebras of binary relations.
@@ 3 Constructing finite odd Sugihara chains
To prove the left-to-right inclusion, assume $x^{b}\leqslant_{X}y^{d}$. Then $x^{b}=y^{d}$ or $x<y$. If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$. Hence, $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, i.e., $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}\circ\alpha=0$. On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$. Again we get $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}=0$.
For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$. Then we have $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, which means $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$. We therefore get $y^{-d}\not\leqslant x^{b}$. If $b=-d$, then it must be the case that $y\not\leqslant x$, in which case we have $x<y$. Hence, $x^{b}\leqslant_{X}y^{d}$. On the other hand, if $b\neq-d$, then $y^{-d}\not\leqslant x^{b}$ implies $y\not<x$. This means $x\leqslant y$, and so, since $d=b$, we get $x^{b}\leqslant_{X}y^{d}$.
@@
Now set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$.
The lemma below will help to simplify some of the proofs later on.
@@
Let $n\geqslant 1$ and $j,k\in\left\{1,\ldots,n\right\}$. If $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$, then $\left(p,r\right)\in L_{\textnormal{min}\{j,k\}}$.
Proof 3.4
Let $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$. Then $\left(p_{1},\ldots,p_{j-1}\right)=\left(q_{1},\ldots,q_{j-1}\right)$, $p_{j}<_{\mathbb{Q}}q_{j}$, $\left(q_{1},\ldots,q_{k-1}\right)=\left(r_{1},\ldots,r_{k-1}\right)$ and $q_{k}<_{\mathbb{Q}}r_{k}$. Clearly then $\left(p_{1},\ldots,p_{\textnormal{min}\{j,k\}-1}\right)=\left(r_{1},\ldots,r_{\textnormal{min}\{j,k\}-1}\right)$. If $j=k$, then we get $p_{j}<q_{j}<r_{j}$. If $\text{min}\{j,k\}=j$, then $p_{j}<q_{j}=r_{j}$. Finally, if $\text{min}\{j,k\}=k$, then $p_{k}=p_{k}=q_{k}<r_{k}$. Hence, in all three cases we have $(p,r)\in L_{\text{min}\{j,k\}}$.
Lemma 3.5
Let $n\geqslant 1$. Then $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
Proof 3.6
The relation $\leqslant_{n}$ is clearly reflexive and anti-symmetric. For transitivity, assume $p\leqslant_{n}q$ and $q\leqslant_{n}r$. If $p=q$ or $q=r$, then we clearly have $p\leqslant_{n}r$. Assume there are $j,k\in\left\{1,\ldots,n\right\}$ such that $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$. In this case we have $\left(p,r\right)\in L_{{\textnormal{min}\{j,k\}}}$ by Lemma 3.3, which means $p\leqslant_{n}r$.
Finally, we show that $\leqslant_{n}$ is a linear order. If $p=q$, we are done, so suppose $p\neq q$. Then there is a smallest $i\in\left\{1,\ldots,n\right\}$ such that $p_{i}\neq q_{i}$. Since the set of rational numbers is linearly ordered, we have $p_{i}<_{\mathbb{Q}}q_{i}$ or $q_{i}<_{\mathbb{Q}}p_{i}$. If $i=1$, then $\left(p,q\right)\in L_{1}$ or $\left(q,p\right)\in L_{1}$, which means $p\leqslant_{n}q$ or $q\leqslant_{n}p$. On the other hand, if $i\in\{2,\ldots,n\}$, then $p_{j}=q_{j}$ for all $j\in\left\{1,\ldots,i-1\right\}$. Hence, $\left(p,q\right)\in L_{i}$ or $\left(q,p\right)\in L_{i}$, and therefore $p\leqslant_{n}q$ or $q\leqslant_{n}p$.
@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
Consider an index set $I$ and let $\mathcal{F}$ be an ultrafilter on $I$. For each index $i\in I$, let $\mathbf{X}_{i}=\left\langle X_{i},\leqslant_{i}\right\rangle$ be a poset and $E_{i}$ an equivalence relation on $X_{i}$ such that $X_{i}\neq\varnothing$ and ${\leqslant_{i}}\subseteq E_{i}$. Consider the Cartesian product $X=\prod_{i\in I}X_{i}$ and the equaliser $\llbracket x=y\rrbracket=\left\{i\in I\mid x(i)=y(i)\right\}$ of $x,y\in X$. Define a binary relation on $X$ by setting, for all $x,y\in X$,
@@
We first prove (i). Let $[x]\in Y$. Since $\leqslant_{i}$ is reflexive for all $i\in I$, we get $\left\{i\in I\mid x(i)\leqslant_{i}x(i)\right\}=I$. Now $\mathcal{F}$ is an ultrafilter, so $I\in\mathcal{F}$. Hence, $\left[x\right]\leqslant_{Y}\left[x\right]$.
Next, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[x]$. Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$. Since $\mathcal{F}$ is closed under taking intersections, it follows that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}.$
Using the fact that $\leqslant_{i}$ is anti-symmetric for all $i\in I$, we can show that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}=\left\{i\in I\mid x(i)=y(i)\right\}.$
Hence we get $\left\{i\in I\mid x(i)=y(i)\right\}\in\mathcal{F}$. This shows that $(x,y)\in E_{\mathcal{F}}$, and hence $[x]=[y]$.
For the transitivity, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[z]$. Then $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$, and so $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$. We know that $\leqslant_{i}$ is transitive for all $i\in I$, so it follows that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}z(i)\right\}\subseteq\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}.$
Hence, since $\mathcal{F}$ is closed under taking supersets, we have $\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}\in\mathcal{F}$, which means $[x]\leqslant_{Y}[z]$.
We next prove item (ii). Firstly, since $E_{i}$ is reflexive for all $i\in I$ and $I\in\mathcal{F}$, it follows that $\left([x],[x]\right)\in E_{Y}$ for all $[x]\in Y$. Using a similar argument as for item (i), we can show that $E_{Y}$ is transitive. For symmetry, assume $\left([x],[y]\right)\in E_{Y}$. Then $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. Since $E_{i}$ is symmetric for all $i\in I$, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}=\left\{i\in I\mid\left(y(i),x(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
To see that ${\leqslant_{Y}}\subseteq E_{Y}$, let $[x]\leqslant_{Y}[y]$. Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$. Since ${\leqslant_{i}}\subseteq E_{i}$ for all $i\in I$, it follows that $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and so $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. Hence, $\left([x],[y]\right)\in E_{Y}$.
The first part of item (iii) is proved as follows:
$[x]\leqslant_{Y}[y]$ iff $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha_{X}\left(x\right)\right)(i)\leqslant_{i}\left(\alpha_{X}\left(y\right)\right)(i)\right\}\in\mathcal{F}$
iff $\left[\alpha_{X}(x)\right]\leqslant_{Y}\left[\alpha_{X}(y)\right]$
iff $\alpha_{Y}\left([x]\right)\leqslant_{Y}\alpha_{Y}\left([y]\right).$
The second equivalence above follows from the equality
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}=\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\},$
which is a consequence of the fact that $\alpha_{i}$ is an order automorphism of $\mathbf{X}_{i}$ for all $i\in I$.
To show that $\alpha_{Y}\subseteq E_{Y}$, let $\left([x],[y]\right)\in\alpha_{Y}$. Then $\alpha_{Y}\left([x]\right)=[y]$, and so $\left[\alpha_{X}(x)\right]=[y]$. Hence, $\left(\alpha_{X}(x),y\right)\in E_{\mathcal{F}}$, and so $\left\{i\in I\mid\left(\alpha_{X}(x)\right)(i)=y(i)\right\}=\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\in\mathcal{F}$. Since $\alpha_{i}\subseteq E_{i}$ for all $i\in I$, we have $\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and therefore, since $\mathcal{F}$ is closed under taking supersets, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. We thus get $\left([x],[y]\right)\in E_{Y}$.
Item (iv) follows from part (i) of Theorem 2.7.
@@
Before proving that the class $\mathsf{RDInFL}$ is closed under ultraproducts, we need the lemmas below to ensure that we have a candidate for the injective homomorphism needed in Theorem 4.11. For Lemmas 4.7 and 4.9, and Theorem 4.11, we will be working in the following setting. Let $I$ be an index set and $\mathcal{F}$ an ultrafilter on $I$. For each $i\in I$, let $\mathbf{A}_{i}$ be a representable distributive InFL-algebra. Let $\mathbf{A}$ denote the direct product $\prod\left\{\mathbf{A}_{i}\mid i\in I\right\}$, and $\mathbf{A}/\theta_{\mathcal{F}}$ the ultraproduct.
@@
$\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}.$
If we can show that $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$ is contained in $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}$, we can use the fact that $\mathcal{F}$ is closed under taking supersets to conclude that $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$, which means $\left([u],[v]\right)\in R_{[a]}$. To this end, let $j\in\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$. Then $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$, $u(j)\leqslant_{j}x(j)$ and $y(j)\leqslant_{j}v(j)$. From the second and third part we obtain $\left(x(j),y(j)\right)\preccurlyeq_{j}\left(u(j),v(j)\right)$. Hence, since $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$ and $\varphi_{j}\left(a(j)\right)$ is an upset of $\left\langle E_{j},\preccurlyeq_{j}\right\rangle$, we have $\left(u(j),v(j)\right)\in\varphi_{j}\left(a(j)\right)$.
@@
Using the facts that $\mathcal{F}$ is an ultrafilter and $\varphi_{i}$ preserves joins for all $i\in I$, we can also show that $\varphi$ preserves joins:
$\left([x],[y]\right)\in\varphi\left([a]\vee^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)$
iff $\left([x],[y]\right)\in\varphi\left(\left[a\vee^{\mathbf{A}}b\right]\right)$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\vee^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\vee^{\mathbf{A}_{i}}b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\cup\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\textnormal{ or }\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cup\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}\textnormal{ or }\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\textnormal{ or }\left([x],[y]\right)\in\varphi\left([b]\right)$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\cup\varphi\left([b]\right).$
@@
We now prove that $\varphi$ preserves --:
$\left([x],[y]\right)\in\varphi\left(-^{\mathbf{A}/\theta_{\mathcal{F}}}[a]\right)=\varphi\left(\left[-^{\mathbf{A}}a\right]\right)$ iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(-^{\mathbf{A}}a\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(-^{\mathbf{A}_{i}}a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in-\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\alpha_{i}\circ\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha_{i}\left(x(i)\right),y(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}.$
We recall that the complement $\left(\varphi_{i}\left(a(i)\right)\right)^{c}$ is being taken in $E_{i}$, hence
$\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\textnormal{ and }\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\cap\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}^{c}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left([y],[\alpha_{X}(x)]\right)\in E_{Y}\textnormal{ and }\left([y],[\alpha_{X}(x)]\right)\notin\varphi\left([a]\right)$
iff $\left([y],[\alpha_{X}(x)]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
iff $\left([\alpha_{X}(x)],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left(\alpha_{Y}\left([x]\right),[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left([x],[y]\right)\in\alpha_{Y}\circ\left(\varphi\left([a]\right)\right)^{c\smile}=-\varphi\left([a]\right).$

### v3

@@ Abstract
We show that all Sugihara monoids can be represented as algebras of binary relations, with the monoid operation given by relational composition. Moreover, the binary relations are weakening relations. The first step is to obtain an explicit relational representation of all finite odd Sugihara chains. Our construction mimics that of Maddux [16], where a relational representation of the finite even Sugihara chains is given. We define the class of representable Sugihara monoids as those which can be represented as so-called direct reducts of distributive involutive FL-algebras of binary relations. We then show that the class of representable distributive involutive FL-algebras is closed under ultraproducts. This fact is used to demonstrate that the two infinite Sugihara monoids that generate the quasivariety are also representable. From this it follows that all Sugihara monoids are representable.
@@ 1 Introduction
For the odd Sugihara chain $\mathbf{S}_{2n+3}$ (where $n\in\mathbb{Z}^{+}$), we will take two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$, with a map $\alpha$ swapping the occurrence of each element from one copy to the other. This map $\alpha$ is an order isomorphism of the poset consisting of two chains. The algebra $\mathbf{S}_{2n+3}$ is then embedded in the upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$, where this time $\mathbf{X}$ consists of two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$ with order relations between them (see Figure 1). The key step is that $\alpha$ is used to define the order-reversing involution in a way that ensures the algebra will be an odd Sugihara monoid.
@@
We will view Sugihara monoids as examples of distributive involutive FL-algebras (DInFL-algebras). See Section 2.1 for a precise definition of these algebras. DInFL-algebras have two order-reversing operations $\sim$ and --. If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic. The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7). Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
@@ 2.1 Sugihara monoids and DInFL-algebras
If $n$ is odd then the monoid identity is given by $1=a_{0}$, and if $n$ is even, then the monoid identity is $1=a_{1}$. This gives us algebras $\mathbf{S}_{n}=\langle S_{n},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$. (When we talk of a Sugihara chain, we mean a linearly ordered Sugihara monoid, using slightly different terminology from that used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
@@ 2.3 Constructing InFL-algebras of relations
Before we give some examples, we need the definition of a direct reduct. This is an adaptation of [14, Definition 8]. It differs from an ordinary reduct because the set of operations is not simply a subset of the original, but new operations are also defined in terms of the old operations.
@@
The two-element Sugihara chain $\mathbf{S}_{2}$ can be represented over $X=\{x\}$ with ${\leqslant}=E=\alpha=\mathrm{id}_{X}$.
@@ 2.4 Representable DInFL-algebras
The following theorem relies on part (iii) of Theorem 2.12 and is used to define representable DInFL-algebras.
@@
[3, Theorem 4.4] $\mathbb{P}\left(\mathsf{FDInFL}\right)=\mathsf{EDInFL}$.
@@
We denote the class of representable DInFL-algebras by $\mathsf{RDInFL}$. In Theorem 4.12 we will show that $\mathsf{RDInFL}$ is closed under ultraproducts. That will be a key step in showing that all Sugihara monoids can be represented as algebras of binary relations.
@@ 3 Constructing finite odd Sugihara chains
To prove the left-to-right inclusion, assume $x^{b}\leqslant_{X}y^{d}$. Then $x^{b}=y^{d}$ or $x<y$. If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{d}\right)\in 0$. On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$. Again we have $(x^{b},y^{d})\in 0$. For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$. Then we have $y^{-d}\not\leqslant x^{b}$. If $b=-d$, then $y\not\leqslant x$, so $x<y$. Hence, $x^{b}\leqslant_{X}y^{d}$. On the other hand, if $b\neq-d$, then $y^{-d}\not\leqslant x^{b}$ implies $y\not<x$. This means $x\leqslant y$, and so, since $d=b$, we get $x^{b}\leqslant_{X}y^{d}$.
@@
Now set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$. Then $\leqslant_{n}$ is the lexicographic order on $\mathbb{Q}^{n}$. Since the rational numbers is linearly ordered, it follows that $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain. Part (i) of the lemma below can be used to show that $\leqslant_{n}$ is transitive. It will also help to simplify some of the proofs later on.
@@
Let $n\geqslant 1$.
  1. (i)
    If $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$ with $j,k\in\{1,\ldots,n\}$, then $\left(p,r\right)\in L_{\textnormal{min}\{j,k\}}$.
  2. (ii)
    The structure $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
Below we outline how to construct the ultraproduct of a set of partially ordered sets with additional structure, each of which can be used to represent a DInFL-algebra. Consider an index set $I$ and let $\mathcal{F}$ be an ultrafilter on $I$. For each index $i\in I$, let $\mathbf{X}_{i}=\left\langle X_{i},\leqslant_{i}\right\rangle$ be a poset and $E_{i}$ an equivalence relation on $X_{i}$ such that $X_{i}\neq\varnothing$ and ${\leqslant_{i}}\subseteq E_{i}$. Consider the Cartesian product $X=\prod_{i\in I}X_{i}$ and the equaliser $\llbracket x=y\rrbracket=\left\{i\in I\mid x(i)=y(i)\right\}$ of $x,y\in X$. Define a binary relation on $X$ by setting, for all $x,y\in X$,
@@
Since all the conditions in items (i) to (iii) can be translated into first-order sentences in the relevant language, they follow from Łoś’s Theorem. Item (iv) follows from part (i) of Theorem 2.7.
Remark 4.7
The fact that the structures with which the DInFL-algebras are represented (i.e. posets with an equivalence relation and an order automorphism) are first-order structures, greatly simplifies the proof that $\mathsf{RDInFL}$ is closed under ultraproducts. Theorem 4.5 above gives us the candidate structure from which we can build the algebra in which we will embed the ultraproduct of representable algebras.
@@
Before proving that the class $\mathsf{RDInFL}$ is closed under ultraproducts, we need the lemmas below to ensure that we have a candidate for the injective homomorphism needed in Theorem 4.12. For Lemmas 4.8 and 4.10, and Theorem 4.12, we will be working in the following setting. Let $I$ be an index set and $\mathcal{F}$ an ultrafilter on $I$. For each $i\in I$, let $\mathbf{A}_{i}$ be a representable distributive InFL-algebra. Let $\mathbf{A}$ denote the direct product $\prod\left\{\mathbf{A}_{i}\mid i\in I\right\}$, and $\mathbf{A}/\theta_{\mathcal{F}}$ the ultraproduct.
@@
$J=\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}.$
If we can show that $J\subseteq\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}$, we can use the fact that $\mathcal{F}$ is closed under taking supersets to conclude that $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$, which means $\left([u],[v]\right)\in R_{[a]}$. To this end, let $j\in J$. Then $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$, $u(j)\leqslant_{j}x(j)$ and $y(j)\leqslant_{j}v(j)$. From the second and third part we obtain $\left(x(j),y(j)\right)\preccurlyeq_{j}\left(u(j),v(j)\right)$. Hence, since $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$ and $\varphi_{j}\left(a(j)\right)$ is an upset of $\left\langle E_{j},\preccurlyeq_{j}\right\rangle$, we have $\left(u(j),v(j)\right)\in\varphi_{j}\left(a(j)\right)$.
@@
Using the facts that $\mathcal{F}$ is an ultrafilter and $\varphi_{i}$ preserves joins for all $i\in I$, we can also show that $\varphi$ preserves joins using a proof dual to the case for meets.